Please help me solve this problem
Jay's mother was 25 years old when he was born. Thirteen years from now, his mother would be twice as old as jay. What are their present ages?

Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

Jay's mother was 25 years old when he was born. Thirteen years from now, his mother would be twice as old as jay. What are their present ages?
----
Equations:
M = J + 25
M+13 = 2(J+13)
--------------
Substitute for "M" and solve for "J"::
J+25+13 = 2J + 26
J = 12 yrs
---
M = J+25 = 37 yrs
-------------
Cheers,
Stan H.
-----

Answer by greenestamps(13200)



The difference in their ages will always be 25 years. So when she is twice as old as her son, she will be 50 and he will be 25.

If that will happen 13 years from now, then their ages now are 50-13=37 and 25-13=12.